OpenStudy (anonymous):

Yolanda received a $80 gift card for a coffee store. She used it in buying some coffee that cost $7.41 per pound. After buying the coffee, she had $57.77 left on her card. How many pounds of coffee did she buy?

OpenStudy (anonymous):

let x=cost per pound, so x*7.41= 80-57.77 x=(80-57.77)/7.41 = 3